Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day13


Crew 100B Log Book for March 10, 2011

Photos of the day





Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Excellent

Experiment Status: All the medical , simulated micro-gravity and human factors experiments were completed successfully .

Time Departed/Returned from EVA: 3 EVAs performed

The meals today were OK since it was a non-cooking day.

EVA’s: We completed three successful EVA.

Number Of reports: one science reports, 1 engineering report, 1 journalist report, 1 EVA report and 1 chef report.

Plan for tomorrow: We will summarize our experiments. Tomorrow will be a cooking day.


Commander's Report

Balwant Rai Reporting

Today we started the day with medical experiments, human factors and our habitability experiment which were completed today.

Today is a non-cooking day. Our crew experts began with analysis of data for their experiments. We successfully completed three EVAs.

Human machine, simulated microgravity experiments, Medical, habitability and astrobiological experiments will be analyzed.


Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): full

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.31

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 1 container full

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 20

· Trailer Potable Water Tank Level (inches from bottom): 2/3

· Hab Potable Water Tank Level (inches from bottom): 7

· Potable Water Meter Reading: 433329


Notes/Comments :


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green plants and snails

· Condition of plants in Tank 3: green plants

· Greenhab notes/ comments:


Observatory:

· Telescope Used? (Yes/no) Yes

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: ok

· Comments:


ATV 2 (Viking I):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: ok

· Comments:


ATV 3 (Viking II):

· Used: Yes

· Oil Checked: no

· Fuel Consumed: Replaced

· Tire Status: ok

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 68

· Actual temperature upstairs (Farenheit): 72

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 62


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:



Science Report

Balwant Rai and Jasdeep Kaur Reporting

Saliva as diagnostic tool in Mars mission

Saliva samples were taken from all crew members (via Versi Sal 1 from Oasis Diagnostics) and preserved for further analysis (Stress markers, inflammatory markers and immunological markers). Salivary amylase and CRP were analyzed by bio-sensor and their levels were significantly higher.


EVA 37 Report

Jasdeep Kaur Reporting

EVA # 37, 14:45-16:00

Crew Members: Quentin Bourges, Jasdeep Kaur, Balwant Rai

Transit Mode: ATVs

Objectives: To build up a well co-ordinated rescue procedure for any crew member in medical distress condition(ENACTED by crew members).

Results: 3 crew members were selected. One crew member enacted to be in a shock condition(enacted by Jasdeep Kaur ), emergency physician started all emergency safety procedures and rescued the injured crew member carrying her on a stature to the Hab where further medical procedures were performed.

Lessons Learned: As per medical report


EVA 38 Report

Matthieu Ansart Reporting

EVA # 37, 14:45-16:00

Crew Members: Quentin Bourges, Matthieu Ansart, Rachel Dompnier

Transit Mode: ATVs

Objectives: Exploration of Mars

Results: We saw cows on the road and we tried not to disturb them.

Lessons learned : We are not alone in the desert. We have to be careful with our environment.


EVA 39 Report

Quentin Bourges Reporting

EVA#39: 5:40 pm to 6:20 pm

Crew members: Balwant Rai, Crystal Latham, Quentin Bourges

Location: 2 km west of the Hab

Transit mode: ATVs

Objective: Exploration of planet Mars

Results: We increased the knowledge of the area around the Hab, following the road.


Chef's Report

Rachel Dompnier Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Rice with vegetables and burger.

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Leftovers from rice and vegetables because I made too much.

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Nothing

5) Lessons learned (if any)

Nothing

6) Comments/questions for Kim and Jean, the food study investigators

Nothing


Journalist Report

Crystal Latham Reporting

Crew 100B has been working hard during their last few days on Mars. Today was a day full of getting results and making conclusions. All of crew 100B diligently worked on finishing up their research projects today.

Today was a non-cooking day, but we had a very delicious lunch none the less. After an interesting turn of events yesterday, we are again trying to make some lime Jell-O. The crew commander thoroughly enjoyed the frozen Jell-O from yesterday; hopefully it turns out more jiggly tonight.

Today members of the crew went on EVA’s to explore the Martian surface more thoroughly and to take a special video. It seems cows are indigenous to both Earth and Mars, since there were a number of sighting while on EVA of the curiously calm creatures. Maybe we Marsonauts aren’t as alone as we thought we were…